# Rigorously Characterizing Dynamics with Machine Learning

### v1

@@ 4.3. Piecewise linear approximations
6. Concluding Remarks
The focus of this paper is on demonstrating via Theorem 1.4 that, at least conceptually, it is possible to characterize global nonlinear dynamics from time series via machine learning. The fundamental idea of Theorem 1.4 is as follows: given any Morse representation ${\mathsf{MR}}$ of the dynamics generated by a continuous map $f$ and a sufficiently good approximation $G$ of $f$, one can construct a combinatorial representation ${\mathcal{G}}$ of the dynamics from which it is possible to reconstruct ${\mathsf{MR}}$.
Of course, in applications $f$ is unknown and hence ${\mathsf{MR}}$ is unknown. Furthermore, the (nontrivial) development of strategies for the machine learning of $G$ and efficient computation of $\mathcal{G}$ are left to future work (a preliminary experimental implementation of a related machine-learning framework is carried out in the context of differential equations in [11]).
Nevertheless, it is possible to discuss how the results of this paper could be used in practice. For the remainder of these concluding remarks we assume that $X\subset{\mathbb{R}}^{d}$ is a finite union of closed, bounded, convex $d$-dimensional polytopes; that $f$ is a continuous function such that $f(X)\subset\mathop{\mathrm{int}}\nolimits(X)$; that $G\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is a fully-connected feedforward neural network with piecewise linear activation; that ${\mathcal{X}}$ is the finite simplicial complex arising from $G$ as defined at the beginning of Section 4.3; and ${\mathcal{G}}_{\rho}$ is the $\rho$-minimal multivalued outer approximation of $G$ on ${\mathcal{X}}$.
We have taken $G$ to be given, but presumably it has been learned based on a finite set of data of the form ${\mathcal{D}}=\left\{{(x,f(x))}\,\left|\,{x\in{\mathbb{R}}^{d}}\right.\right\}$.
As detailed in Section 5 and stated in Theorem 1.4, for each ${\mathcal{G}}_{\rho}$ we construct a Morse graph ${\mathsf{MG}}({\mathcal{G}}_{\rho})$ for which there exists a poset epimorphism $\nu_{\rho}\colon{\mathsf{MG}}({\mathcal{G}}_{\rho})\to{\mathsf{MR}}$, and $\nu_{\rho}$ can be used to determined $\mathop{\mathrm{Con}_{*}}\nolimits(M)$ for $M\in{\mathsf{MR}}$. Of course, since ${\mathsf{MR}}$ is not known, $\nu_{\rho}$ is not a priori determined. However, if we choose $\nu_{\rho}$ to be an poset epimorphism, then $\nu_{\rho}({\mathsf{MG}}({\mathcal{G}}_{\rho}))$ defines a Morse representation of $f$.
The Conley index may be used to deduce the existence of a nontrivial Morse representation for $f$. To demonstrate this, we return to the ideas of Remark 2.18. Consider a finite sublattice ${\mathsf{N}}$ of ${\mathsf{ABlock}}(f)$ and assume that $k\colon{\mathsf{A}}\to{\mathsf{N}}$ is a lattice monomorphism where ${\mathsf{A}}$ is a finite sublattice of ${\mathsf{Att}}(f)$. [^1] Note that the trivial lattice monomorphism always exists; choose ${\mathsf{A}}=\left\{{\emptyset,\omega(X,f)}\right\}$ and set $k(\emptyset)={\mathbf{0}}_{b}$ and $k(\omega(X,f))={\mathbf{1}}_{b}$. In this case, the associated Morse representation, ${\mathsf{MR}}({\mathsf{A}})=\left\{{\omega(X,f)}\right\}$, is trivial. Assume that there exists $N_{1},N_{0}\in{\mathsf{J}}^{\vee}({\mathsf{N}})$ such that $N_{0}\subset N_{1}$ and $N_{0}\neq{\mathbf{0}}_{b}$. If $f_{*}\colon H_{*}(N_{1},N_{0})\to H_{*}(N_{1},N_{0})$ is not shift equivalent to the trivial map, then $M=\mathop{\mathrm{Inv}}\nolimits(N_{1}\setminus N_{0},f)\neq\emptyset$ and thus we can conclude that there exists a nontrivial Morse representation.
It is straightforward to check that sufficiently large outer approximations of $G$ define an outer approximation of $f$, but due to its importance we state the result as a proposition.
Proposition 6.1.
Assume that ${\mathcal{G}}_{\rho_{0}}$ is an outer approximation of $f$ on ${\mathcal{X}}$ for some fixed $\rho_{0}>0$. Then, for all $\rho\geq\rho_{0}$, ${\mathcal{G}}_{\rho}$ is an outer approximation of $f$ on ${\mathcal{X}}$.
Without loss of generality assume that $\mathop{\mathrm{diam}}\nolimits(X)=1$. In this case ${\mathcal{G}}_{1}(\xi)={\mathcal{X}}$ for each $\xi\in{\mathcal{X}}$. The Morse graph ${\mathsf{MG}}({\mathcal{G}}_{1})$ consists of a single node. Observe that this does provide a Morse representation of $f$, in other words, we have truth but no useful information. Since $G$ and ${\mathcal{X}}$ are assumed to be fixed and ${\mathcal{X}}$ is finite, there are a finite set of combinatorial multivalued maps ${\mathcal{G}}_{\rho}\colon{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ as $\rho$ ranges over the interval $[0,1]$.
Let $\rho_{0}$ be the minimal value for which ${\mathcal{G}}_{\rho_{0}}$ is an outer approximation of $f$ on ${\mathcal{X}}$ ($\rho_{0}$ is unknown). For $\rho\in[0,\rho_{0})$, the Morse graph ${\mathsf{MG}}({\mathcal{G}}_{\rho})$ does not necessarily provide correct information about the dynamics of $f$. However, for $\rho\in[\rho_{0},1]$ the Morse graph ${\mathsf{MG}}({\mathcal{G}}_{\rho})$ provides correct information about the dynamics of $f$. In principle, collecting more data (enlarging the set of data ${\mathcal{D}}$) would result in $G^{\prime}$ that provides a better approximation of $f$, a finer simplicial complex ${\mathcal{X}}^{\prime}$, and a smaller $\rho^{\prime}_{0}$ for which ${\mathcal{G}}^{\prime}_{\rho^{\prime}_{0}}$ provides an outer approximation of $f$ on ${\mathcal{X}}^{\prime}$. Correct information about the dynamics remains correct as more data is collected, which suggests that information about dynamics that persists as more data is collected may be trustworthy. It is conceivable that this last statement could be quantified given Lipschitz bounds for $f$ over $X$, which we leave for future research.
Acknowledgements
B.G. was supported by the National Science Foundation under Grant DGE-1842213. M.G. and K.M. were partially supported by the Air Force Office of Scientific Research under awards numbered FA9550-23-1-0011 and FA9550-23-1-0400.
7. Appendix
We recall elementary definitions associated with lattices and refer readers to [6] for a more comprehensive introduction.
Definition 7.1.
A lattice is a set ${\mathsf{L}}$ together with binary operations $\wedge$ and $\vee$ that are commutative and associative and furthermore satisfy the following properties.
  1. (i)
    $(\forall a\in{\mathsf{L}})$ $a\vee a=a=a\wedge a$.
  2. (ii)
    $(\forall a,b\in{\mathsf{L}})$ $a\wedge(a\vee b)=a\vee(a\wedge b)=a$.
A lattice is distributive if it satisfies the distributive law,
$(\forall a,b,c\in{\mathsf{L}})\hskip 5.69046pta\wedge(b\vee c)=(a\wedge b)\vee(a\wedge c).$
The meet operation induces a partial order defined by $a\leq_{\mathsf{L}}b$ if and only if $a\wedge b=a$.
A lattice is bounded if there exist minimal and maximal elements ${\mathbf{0}}$ and ${\mathbf{1}}$ with respect to the partial order $\leq_{\mathsf{L}}$. Throughout this paper we always work with bounded lattices.
A lattice morphism $f\colon{\mathsf{L}}\to{\mathsf{L}}^{\prime}$ between two bounded lattices ${\mathsf{L}}$ and ${\mathsf{L}}^{\prime}$ preserves the lattice operations, i.e., $f(a\vee b)=f(a)\vee f(b)$ and $f(a\wedge b)=f(a)\wedge f(b)$ for all $a,b\in{\mathsf{L}}$, and $f({\mathbf{0}})$ and $f({\mathbf{1}})$ are minimal and maximal elements in ${\mathsf{L}}^{\prime}$.
An element $c\in{\mathsf{L}}$ is join-irreducible if there exists a unique element $c^{<}\neq c\in L$ such that for all $b<c$, $b\leq c^{<}$. In this case, $c^{<}$ is called the immediate predecessor of $c$. The set of join-irreducible elements of a lattice $L$ is denoted by ${\mathsf{J}}^{\vee}({\mathsf{L}})$. The predecessor map :${}^{<}:{\mathsf{J}}^{\vee}({\mathsf{L}})\to L$ sends a join-irreducible element $c\in{\mathsf{L}}$ to its unique predecessor, $c^{<}$.

### v2

@@ 4.3. Approximating attracting blocks
6. Examples
In this section, to discuss the applicability of Theorem 1.2, we generate synthetic data using a two-dimensional nonlinear population model,
(5) $f_{\theta}(x):=\begin{bmatrix}(\theta_{1}x_{1}+\theta_{2}x_{2})e^{-0.1(x_{1}+x_{2})}\\ 0.7x_{1}\end{bmatrix}$
where the parameter $\theta$ is chosen to be ${\bar{\theta}}=(23.5,23.5)$. As this is meant to be a population model it makes sense to restrict our attention to data in the non-negative orthant $0,\infty)^{2}$ and we leave it to the reader to check that for $X=[0,90]\times[0,70]$, $f_{\bar{\theta}}(X)\subset\mathop{\mathrm{int}}\nolimits(X)$.
We choose (5) for four reasons. First, mathematically rigorous computations of the dynamics generated by $f_{\theta}$ can be performed (see [arai:kalies:kokubu:mischaikow:oka:pilarczyk, bush:gameiro:harker:kokubu:mischaikow:obayashi:pilarczyk]). Second, it is known that at $\bar{\theta}$}, the dynamics is complicated (see [ugarcovici:weiss, Figure 1]). Third, as a function of $\theta$ the dynamics of $f_{\theta}$ undergoes dramatic changes over complicated sets of bifurcation points (see [ugarcovici:weiss, Section 5]). Fourth, we can demonstrate how, with relatively few data points, we can capture interesting dynamics, and indicate where failure is to be expected.
6.1. Baseline
To obtain a baseline, we use the software package [CMGDB] to construct multivalued maps by evaluating $f_{\bar{\theta}}$ on the vertices of a cubical cell complex obtained by uniformly subdividing $X=[0,90]\times[0,70]$. We consider two levels of subdivision, and therefore two distinct cell complexes – ${\mathcal{Z}}_{1}$ and ${\mathcal{Z}}_{2}$ consisting of $2^{18}$ and $2^{27}$ rectangles, respectively – and corresponding multivalued maps ${\mathcal{F}}_{1}$ and ${\mathcal{F}}_{2}$. Figure 4 displays the Hasse diagrams of the Morse graphs together with the Conley indices and the corresponding color-coded subsets of phase space. Comparing Figures 4(a) and (b), we observe that the greater level of subdivision leads to a more refined description of the dynamics.
Each node $q_{k}$ of the Morse graph has a label of the form $k:(s_{0}(x),s_{1}(x),s_{2}(x))$ where $k$ identifies the node and $s_{i}(x)$ is the product of the invariant factors (excluding the nilpotent blocks) [hoffman:kunze, Section 7.2] of the $i$-th homology map of the index pair (see Definition 2.9). Since $X\subset{\mathbb{R}}^{2}$, $s_{i}(x)\equiv 0$ for $i\geq 3$. We use ${\mathbb{F}}_{5}$ coefficients for the computations presented here.
Turning to Figure 4(a), the Morse graph ${\mathsf{MG}}({\mathcal{F}}_{1})$ has four nodes $\left\{{0,1,2,3}\right\}$. The Conley index of node 0, $(x^{3}-1,0,0)$, is that of a stable period three orbit. From this we can conclude that $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(0)|,f_{\bar{\theta}})$ contains a period three orbit (though it need not be stable). Similarly, the Conley index of node $1$ is that of an unstable (orientation preserving) period-3 orbit and hence $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(1)|,f_{\bar{\theta}})$ contains a period three orbit. Finally, the Conley index of node $3$ is that of a two-dimensional unstable (orientation preserving) fixed point and hence $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(3)|,f_{\bar{\theta}})$ contains a fixed point. Note that in each case we obtained lower bounds on the dynamics of the invariant sets associated with the Morse nodes. In fact, the dynamics is much more complicated and as indicated in the introduction this section, for some values of $\theta$, too complicated to be completely characterized by finite data (see [ugarcovici:weiss, Section 5]), while our characterization via Morse graphs and Conley indices can be made rigorous (see [arai:kalies:kokubu:mischaikow:oka:pilarczyk]).
The Conley index of node 2 is trivial. However, $|\pi^{-1}(2)|$ contains the origin, which is an unstable fixed point for (5). Thus, $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(2)|,f_{\bar{\theta}})\neq\emptyset$.
Figure 4. Combinatorial dynamics of $\mathcal{F}$} where $\mathcal{F}$} is a multivalued map computed directly using $f_{\bar{\theta}}$ given by (5). We use these results as a baseline for comparison to the results in Figure 5. In both figures, the Hasse diagram of the Morse graph ${\mathsf{MG}}({\mathcal{F}})$ is displayed above a plot of phase space where the sets $|\pi^{-1}(q)|$ are plotted with colors corresponding to $q\in{\mathsf{MG}}({\mathcal{F}})$. The cubical cell complex $\mathcal{X}$} used to obtain these results is constructed by uniformly subdividing phase space into (a) $2^{18}$ rectangles and (b) $2^{27}$ rectangles. For visualization purposes, in plot (a), $|q^{-1}(2)|$ is enlarged by a factor of 10. In plot (b), $|q^{-1}(k)|$ is enlarged by a factor of 100 for $k=2,3,5$, and $|q^{-1}(4)|$ is enlarged by a factor of 200.
We remark that since the invariant sets associated with each node of the Morse graph are nontrivial, we have identified a Morse representation of $f_{\bar{\theta}}$, which we denote by ${\mathsf{MR}}_{1}$. As indicated in the introduction, the use of the language of Morse representations allows us to analyze dynamics at different levels of resolution.
Figure 4(b) provides a finer level of resolution. The Morse graph has six nodes $\left\{{0,1,2,3,4,5}\right\}$. Nodes 0 and $1$ are minimal indicating bistability. The fundamental differences between ${\mathsf{MG}}({\mathcal{F}}_{1})$ and ${\mathsf{MG}}({\mathcal{F}}_{2})$ are that ${\mathcal{F}}_{2}$ reveals that the Morse sets defined by $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(1)|,f_{\bar{\theta}})$ and $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(0)|,f_{\bar{\theta}})$ in Figure 4(a) can each be decomposed into two Morse sets. In particular, in Figure 4(b), $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(0)|,f_{\bar{\theta}})$ has the Conley index of an invariant circle and $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(3)|,f_{\bar{\theta}})$ has the Conley index of an unstable (orientation preserving) period-3 orbit, while $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(1)|,f_{\bar{\theta}})$ has the Conley index of a stable period-6 orbit and $\mathop{\mathrm{Inv}}\nolimits(|\pi^{-1}(2)|,f_{\bar{\theta}})$ has the Conley index of an unstable (orientation reversing) period-3 orbit. We denote the Morse representation associated to this example by ${\mathsf{MR}}_{2}$.
6.2. Model and Training
As proof of concept, we use a standard machine learning model to approximate $f$ and demonstrate Theorem 1.2. We construct data sets of the following form
[…]
where the initial conditions $x_{i}$ are chosen uniformly at random in $X$, and we train using the pairs $\left\{{(f_{\bar{\theta}}^{k-1}(x_{i}),f_{\bar{\theta}}^{k}(x_{i}))}\right\}$.
Fix $G$ to be a fully-connected feedforward neural network with ReLU activation, depth three, and layers of identical width $w$. We train the parameters of $G$ using PyTorch [pytorch] by minimizing the mean squared error with the Adam optimizer [kingma:ba]. The learning rate is initialized at $0.0001$ and reduced on plateau.
Similarly to the baseline method, we uniformly subdivide phase space to obtain a cubical cell complex. Using [CMGDB], we construct a multivalued map $\mathcal{G}$} by evaluating $G$ on the vertices of each rectangle.
6.3. Results
We present a few simple results, but encourage the interested reader to explore with additional experiments [RCDML].
Figure 5. The Hasse diagrams of the Morse graphs computed using neural network approximations to $f_{\theta}$ as defined by Equation 5.
6.3.1. Experiment 1
The results presented in Figure 5(a) were obtained using ${\mathcal{D}}(10,16)$, $w=32$, and $2^{20}$ rectangles. In this case $\nu\colon{\mathsf{MG}}\to{\mathsf{MR}}_{1}$ of Theorem 1.2 is an embedding, but not an epimorphism. This result is easily reproducible: in an experiment with ten trials fixing $w$, $n$, and the number of rectangles, but varying the initial weights of the network at the start of training, we observe the same Morse graph in every trial.
What $\nu$ does not identify is the Morse set associated with the unstable fixed point at the origin. However, this is to be expected. Naively, one expects that for $G$ to learn that there is recurrence near the origin requires the existence of a data point in the rectangle containing the origin whose image is also in that rectangle. A simple computation shows that the value of the unstable eigenvalue at the origin is approximately 24 and thus the area of the region containing such data points is roughly $1.6\times 10^{-7}$ the area of $X$. Therefore, to reliably obtain such a data point suggests the need for on the order of $10^{10}$ data points.
Here we appeal to the aphorism “perfect is the enemy of the good.” Theorem 1.2 can be achieved but at an unreasonable cost compared to the fact that the interesting complex dynamics can be obtained with relatively few data points. It is also worth noting that the difficulty of identifying the fixed point at the origin is due to the fact that it is unstable and lies on the boundary of $X$. The Morse set associated with node 3 in Figure 4(a) is more unstable, but as indicated by node 2 in Figure 1(a) is easily detected.
6.3.2. Experiment 2
In an experiment with ten trials, choosing ${\mathcal{D}}(10,32)$, $w=32$, and $2^{21}$ rectangles, in six trials we computed the Morse graph of Figure 5(b). In the remaining four trials, we again observe the Morse graph of Figure 5(a). The additional spatial refinement allows for the identification of node 0 with Conley index of a stable period-6 orbit. In the case of the refined Morse graph the lattice morphism $\nu\colon{\mathsf{MG}}\to{\mathsf{MR}}_{1}$ of Theorem 1.2 maps nodes 0 and 1 of $\mathsf{MG}$} to the node 0 of ${\mathsf{MR}}_{1}$.
6.3.3. Experiment 3
In an experiment with ten trials, choosing ${\mathcal{D}}(20,128)$, $w=256$, and $2^{23}$ rectangles, we computed the Morse graphs shown in Figure 5(a), (b), (c) and (d), in five, two, one, and two trials, respectively.
As expected the Morse set associated with the unstable fixed point at the origin is still not detected. This example allows us to highlight that the morphism $\nu$ of Theorem 1.2 is dependent on the choice of Morse representation. In particular, $\nu\colon{\mathsf{MG}}\to{\mathsf{MR}}_{1}$ is well defined for all ten trials. However, for the Morse graph shown in Figure 5(d) we obtain $\nu\colon{\mathsf{MG}}\to{\mathsf{MR}}_{2}$ is an embedding. It is worth noting that the embedding is obtained even though the data driven computations were done using $2^{23}$ rectangles as opposed to the $2^{27}$ rectangles used to establish ${\mathsf{MR}}_{2}$.
7. Data and Code Availability
The data and code used to perform the computations are available at [RCDML].